OpenStudy (anonymous):

What is the area of a right triangle with sides of 5, 12, and 13?

OpenStudy (anonymous):

5*12*13= answer

Parth (parthkohli):

The area of a triangle is \(0.5 bh\). Here, \(b = 5\) and \(h = 12\). What is \(0.5 \times 12 \times 5\)?